if it takes 5 quarts to paint a 10m by 30m wall, how many quarts will it take to paint a 30m by 90m wall?
Answer by ewatrrr(24785)

 
Hi,
if it takes 5 quarts to paint a 10m by 30m wall(300m^2 area),
how many quarts will it take to paint a 30m by 90m(2700m^2 area) wall
Let x be the number of qts needed for 2700m^2 wall.
Question sets Up the Proportion
5/300 = x/2700 | Cross Multiplying
300x = 5*2700
x = 5*2700/300
x = 45 qt
CHECKING our Answer***
1/60 = 45/2700